OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations. A) (-2/3, 25/3) B) (2/3, 17/3) C) 2/3, 25/3)

OpenStudy (astrophysics):

Does the method matter?

OpenStudy (anonymous):

no it doe not.

OpenStudy (astrophysics):

Eh, lets use substitution, so for that we need to isolate one of the variables x or y from one of the equations then substitute it into the other equation. So for instance lets take equation 1. 2x+y =7 and solve for y, y = 7-2x We can take this and sub it into the second equation for y, 5x+(7-2x)=9, now solve for x then once you have x sub it back into one of the original equations and solve for y.

OpenStudy (anonymous):

thank you